OpenStudy (anonymous):

A regular octagon is inscribed in a circle of radius r what is the perimeter of the octagon. PLEASE SHARE THE APPROACH

jhonyy9 (jhonyy9):

how do you think what mean ,,octagon" ?

OpenStudy (anonymous):

a polygon with eight sides

jhonyy9 (jhonyy9):

yes,nice so can you subscribing it inside a circle ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

ok but before what mean ,,regular" octagon ?

OpenStudy (anonymous):

regular octagon means all the sides are equal i guess

jhonyy9 (jhonyy9):

yes ,right so than can you draw it please ?

OpenStudy (anonymous):

yes the above pic is perfect

OpenStudy (anonymous):

1080 sum of interieo angles of octagon 135 is the angle between two sides

OpenStudy (anonymous):

but we still need to find that one particualr odd side how to find that

OpenStudy (anonymous):

through angle draw figure such that it form right triangle then tan theta=r/l l=r/tan45

OpenStudy (anonymous):

jhonny bro please take part in ans

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

so how you calcule the opposite side to angle with 45 degree ?

OpenStudy (anonymous):

dont have a clue sir

jhonyy9 (jhonyy9):